OpenStudy (anonymous):

What are the solutions of the system y=x^2 - 3x + 2 ,y=4x -4

OpenStudy (anonymous):

y=x^2 - 3x + 2 -(1) y=4x -4 -(2) Sub (2) into (1) 4x -4 =x^2 - 3x + 2 x^2 - 7x +6 =0 (x-6)(x-1) =0 x =6 or x=1 Now, substitute x =6 and x=1 (2) respectively. When x =6 y=4(6) -4 = 24-4 =20 When x =1 y=4(1) -4 = 4-4 =0 Therefore the solutions are (1,0) and (6, 20) Hope it helps

OpenStudy (anonymous):

Thanks Roly Poly

OpenStudy (anonymous):

Do you understand how to do it?

OpenStudy (amistre64):

a little advice, since you provided the solutions and didnt really engage the asker; we will never really know if they understood ....

OpenStudy (anonymous):

Oh.. Thanks for the advice!